A gold doubloon 6.1 cm in diameter and 2.0mm thick is dropped over the side of a pirate ship. When it comes to rest on the ocean floor at a depth of 770 m, How much has its Volume changed

1
Expert's answer
2021-11-29T11:42:39-0500

Area of the flat surface of the coin:


"A=\pi d^2/4."

Initial volume:


"V=Ah=\frac{\pi d^2h}{4}."

Pressure at 770 m:


"p=\rho gh."

Force on the flat surface of the coin (assuming it is parallel to the ground):


"F=pA=\frac14\rho \pi gh^2d^2."


Stress caused by the force:


"\sigma=\frac FA=p."


Strain in the coin:


"\epsilon=\frac{\Delta h}h=\frac p E,\\\space\\
\Delta h=h\frac pE=\frac{\rho gh^2}{E},"

the volume has shrunk by


"\Delta V=\Delta hA,\\\space\\
\Delta V=\frac{\rho \pi gh^2d^2}{4E}=1.49·10^{-15}\text{ m}^3."

The density of water is 1030 kg, Young's modulus for gold is 79 GPa.